OpenStudy (anonymous):

What is the turning point of the graph of f(x) = |x – 4| ?

OpenStudy (jdoe0001):

|x – 4| <--- what value of "x" makes that expression = 0?

OpenStudy (anonymous):

Ix-4I is x-4

OpenStudy (anonymous):

i dont know.. o.O

OpenStudy (anonymous):

Just to make sure you understand what |....| means... What is |-4| ? What is | 7 | ?

OpenStudy (anonymous):

what's -4? what is 7? i dont get it?

OpenStudy (anonymous):

Those 2 lines mean "absolute value" To take the "absolute value" of a number means simply this: If it is negative, change it to positive. If it is not negative, leave it alone. |-4| = 4 | 7 | = 7

OpenStudy (anonymous):

What do you need to replace x with to make x - 4 be 0 ?

OpenStudy (anonymous):

0?

OpenStudy (jdoe0001):

let us try a few values for "x" see if they work to make it 0 \(\large \begin{array}{llll} x\qquad &f(x)\\ \hline\\ {\color{blue}{ 5}}& |{\color{blue}{ 5}} - 4|\implies |1|\implies 1\\ {\color{blue}{ 2}}&|{\color{blue}{ 2}} - 4|\implies |-2|\implies 2\\ {\color{blue}{ -7}}&|{\color{blue}{ -7}} - 4|\implies |-11|\implies 11\\ \end{array}\) so... those values didn't work so... what do you think will work for "x" to make the expression 0?

OpenStudy (anonymous):

0 x 4?

OpenStudy (jdoe0001):

well, yes if we were multiplying times 4, but we're not we're just subtracting

OpenStudy (anonymous):

x - 4 = 0 x = 4 f(4) = |x - 4| ?

OpenStudy (jdoe0001):

ohhh so if we set x = 4, then the expression turns 0, I see so \(\bf f(x) = |x - 4| \implies f(x) = |x - 4| +0 \\ \quad \\ |{\color{red}{ 4}} - 4| +{\color{red}{ 0}}\\ ({\color{red}{ 4,0}})\Leftarrow \textit{"turning point" or }vertex\)

OpenStudy (anonymous):

thank you all :]

OpenStudy (jdoe0001):

yw